Find the value of 9w + l given that -4w-5 = 3
Simplify your answer as much as possible

Answer by user_dude2008(1862)

-4w-5 = 3
-4w = 3+5
-4w = 8
w = 8/(-4)
w=-2


9w+1=9(-2)+1=-18+1=-17